OpenStudy (anonymous):

Simplify 10 log 2003 without a calculator. Remember that 10 log (2) = 3dB. Not sure which direction to go. I get 10log2003/log10 or 3dB + 10log(2003/2) when I try to use the hint.

OpenStudy (kropot72):

10 log 2003 = 10(log2 + log 1003) The question is suggesting using the value of 0.3 for log 2, which rounds the actual value to 2 decimal places. Therefore we can use 3.0 for log 1000. Can you now find the answer?

OpenStudy (kropot72):

10 log 2003 = 10(log 2 + log 1003) = 10(0.3 + 3.0) = ?

OpenStudy (kropot72):

@optiquest Are you there?

OpenStudy (anonymous):

Sorry was working another problem

OpenStudy (anonymous):

does 10(log2 + log 1003) = 10(log 2006) ?

OpenStudy (kropot72):

Yes you are correct. However if we write: 10(log 2003) = 10(log 2 + log 1001.5) the right hand side would still become 10(0.3 + 3.0) for the same reason i gave previously.

OpenStudy (anonymous):

Oh duh lol

OpenStudy (kropot72):

@optiquest Are you cool with the answer now?

OpenStudy (anonymous):

Yes thank you

OpenStudy (kropot72):

You're welcome :)